# Quantum theory of fractional topological pumping of lattice solitons

### v2

@@ IV Chern number and Wilson loop of solitons in the AAH model
Figure 3: Instantaneous soliton energies for 10 particles for attractive interaction $U=0.1$ with phase-offset $k=1$ in (a) and with phase-offset $k=2$ in (c), and unit-cell size $p=5$. The corresponding COM-movements in a pump cycle of the red and blue marked energies are shown in (b) and (d). The solutions are obtained with the three-site soliton ansatz.
In Fig. 3 we have plotted the energies of the soliton bands for the AAH model, Eq. (1), and the COM transport, for $N=10$ particles with unit cell size $p=5$ and $U=0.1$ for phase offsets $k=1$ and $k=2$. In the first case (Fig. 3 a, b) there is a single lowest-energy soliton band with many-body Chern number, Eq. (9), $C=1$ and integer transport. In the second case (Fig. 3 c, d) two bands cross with a non-trivial Wilson loop, Eq. (16), $C_{2}=1$ giving rise to fractional transport of $1/2$.
@@ V Topological phase transitions
In the following we provide an explanation for the transitions between phases with integer, fractional and vanishing transport observed in the experiments. The COM dynamics of bound $N$-particle objects (solitons) can be described by an effective single-particle Hamiltonian with eigenstates $|E_{\mu}(K)\rangle$:
@@
Since such a model misses out the continuum of extended states, it is only adequate for stable soliton bands. Defining annihilation and creation operators for solitons centered at lattice site $l$ as $\hat{d}_{l}$ and $\hat{d}_{l}^{\dagger}$, respectively, the effective soliton Hamiltonian would read in coordinate space
@@
where the $J_{l,\textrm{eff}}$ are the effective hopping rates of the soliton and the $\epsilon_{l,\textrm{eff}}$ effective local energies. (Note that we here have assumed only nearest neighbor hopping.) All topological properties, including the topological classification according to the Altland-Zirnbauer scheme [24, 25], as well as all phase transitions of solitons are determined by this effective Hamilton. We will explicitly construct $\mathcal{H}_{\textrm{eff}}$ for $N=3$ particles in the strong interaction limit in Sec. VII. Here we will first discuss some of its general properties.
@@
In the Aubry-Andre-Harper model, Eq. (1), any small attraction $U_{0}$ is sufficient to form a bound state (lattice soliton) with an energy band below the continuum of extended $N$-particle states, if $N$ is large. Increasing the attractive interaction, the energy of these bands is lowered and the bands deform. Moreover, excited soliton bands can emerge. If the soliton bands have a non-trival Chern number or Wilson loop, (fractional) quantized topological transport can be observed [6, 13].
@@
The experiments in [6, 13] and DNLSE simulations showed that for very large but still finite values of the interaction, the topological transport stops altogether. This can be understood as follows: In the large $U$ limit the localization length of all soliton solutions is reduced to a single lattice site. Thus the contribution of the local interaction to the energy, UN(N1)/2UN(N-1)/2, is the same for all soliton bands and the total number of stable soliton bands in this limit is equal to the number of single-particle bands. Using a perturbative ansatz one can show that moreover the contribution of the kinetic energy can be disregarded as it scales as $\sim J(\frac{J}{U})^{N-1}$ (see Appendix and Ref. [32]). Thus the effective soliton Hamiltonian becomes approximately diagonal and is entirely determined by local energies $\epsilon_{l,\textrm{eff}}$, which for the AAH model result from virtual hopping processes of a single particle from the soliton site $l$ to an empty neighboring site (and back) and are in second-order perturabtion theory proportional to $\epsilon_{l}\sim(J_{l}^{2}(t)+J_{l+1}^{2}(t))/U$. As a consequence, all soliton bands cross at different points of the pump cycle and the topological transport is given by the total Wilson loop of all COM bands, which is always zero for general reasons.
The two limiting cases show that as $U$ is increased, initially separated soliton bands must approach each other and eventually merge. As shown in Fig. 4, obtained from exact diagonalization (ED) simulations of a small system (30 lattices sites), they do so by forming Dirac-like cones which touch at a critical interaction $U_{c}$. In this case the Chern numbers of the lowest two soliton bands are no longer adequate topological invariants, but one has to consider the Wilson loop. The topological transport can then become fractional, provided the soliton solutions which started at $t=0$ in energetically well separated bands flip bands at the Dirac-like points. We now show that this is indeed the case:
@@
Figure 4: Merging of soliton energies obtained from exact diagonalization with increasing interaction strength $U$ for $N=3$ particles and $p=5$, $k=2$. The number of unit cells is $L=6$. A soliton originally prepared at $t=0$ in one of the two bands will remain in this band if the changes are adiabatic and the soliton bands do not touch (Figs. a) and b)). Once the bands touch in Dirac-like cones, the solution switches bands at every crossing (red curve in Fig. c). The width of the lines reflects the width of the soliton bands in $K$ space.
@@
Close to the crossing points, which we assume to take place at $t=t_{0}$, the non-Abelian Berry phase $\mathbf{A}_{K}$ takes on the form
$\mathbf{A}_{K}=\left(\begin{array}[]{cc}a|\tau|&\frac{ib}{\tau}\\ -\frac{ib}{\tau}&-a|\tau|\end{array}\right),$ (24)
@@
with $\tau=t-t_{0}$. Here, we assumed that apart from a common energy offset $E_{0}=-E_{1}=a\,|\tau|$, and thus $\langle E_{0}|\partial_{t}{\cal H}|E_{1}\rangle/(E_{1}-E_{0})\approx b/\tau$. Due to the diverging off-diagonal elements, the eigensolutions no longer follow the original bands but cross from one band to the other (see color code in Fig. 4 c). Since $p$ is a prime number, there is an odd number of crossing points in one period in the AAH model. Thus, two periods are required for a soliton to return to the band it started from. The shift in the COM position is then integer quantized only after two periods and the average transport per cycle is fractionally quantized.
VI Failure of transport quantization
@@
Based on solutions of the mean-field DNLSE it was shown in Ref. [15] for the example of the Rice-Mele model [33, 34, 35] with local attractive interactions that the transition between quantized topological pumping of a soliton to zero pumping upon increasing the interaction strength may go through an intermediate regime of non-quantized transport. This transient failure of transport quantization, where $\langle\Delta\hat{X}\rangle$ strongly fluctuates, was attributed in [15] to a ”self-crossing” of solutions of the semiclassical DNLSE. This phenomenon can be easily understood in the full quantum picture. For intermediate interactions, the excited soliton band is in some parts of the pump cycle degenerate with the continuum of extended states and thus unstable. If this band merges with the lower soliton band, the time evolution is no longer adiabatic and the transport is not quantized until the interaction is large enough such that also the excited band becomes fully gapped. In Fig. 5 we show the soliton bands in the Rice-Mele model with different values of the attractive onsite interaction $U$ obtained from exact diagonalization simulations of a $N=4$ soliton. Although only marginally visible due to finite size effects, Fig. 5b indicates the existence of a parameter range where the lowest soliton band merges with an unstable excited band leading to a fluctuating, non-quantized transport.
@@
Figure 5: Breakdown of transport quantization in the Rice-Mele model for intermediate values of onsite interactions. Shown are lowest energies for $N=4$ bosons. a) $U=2.0$: lowest soliton (quadron) band is gapped with Chern number $C=1$ ($\langle\Delta\hat{X}\rangle=1$). b) $U=3.0$: lowest two bands merge but (except for finite-size effects) overlap with extended states (unstable quadron) where the transport is not quantized. c) $U=4.0$ lowest two quadron bands merge with vanishing Wilson loop ($\langle\Delta\hat{X}\rangle=0$).
VII Effective Hamiltonian: triplon model
@@
We can explicitly construct the effective soliton Hamiltonian in the strongly interacting limit where the non-interacting part of the Hamiltonian, Eq. (1), can be treated as a small perturbation and the solitons become maximally localized (i.e. with a localization length $\xi$ smaller than the lattice spacing).
@@
We developed a full quantum description of topological pumps of self-bound $N$-particle states, i.e. lattice solitons. The quantum description allowed us to identify and to explicitly calculate a topological invariant, i.e. an effective single-particle Chern-number or Wilson-loop explaining the emergence of integer and fractional transport in the full range of interaction strength, where perturbative arguments [19, 14] fail. Transitions between phases of differently quantized topological transport observed in experiments [6, 13] as well as the possibility of parameter regimes without quantized transport were explained by coalescence of soliton bands and possible degeneracies with extended states. The concept can easily be extended to multi-component solitons [39, 20, 21], were, among other things, fractional transport was predicted despite the fact that all single-particle bands are topologically trivial. In the latter case interactions not only lead to the modification of topological properties, such as the transition from integer to fractional phases, but to the emergence of topology.

### v3

@@ IV Calculating Chern number and Wilson loop of solitons
Figure 3: Instantaneous soliton energies for 10 particles for attractive interaction $U/J=0.1$ with phase-offset $k=1$ in (a) and with phase-offset $k=2$ in (c), and unit-cell size $p=5$. Note that the soliton bands are almost flat and the variation of $E(K)$ with $K$ is less than the width of the lines. The corresponding COM-movements in a pump cycle of the red and blue marked energies are shown in (b) and (d). The energies and COM-positions are obtained with the three-site soliton ansatz.
In Fig. 3 we have plotted the energies of the solitons for the AAH model, Eq. (1), and the COM transport as function of time, for $N=10$ particles with unit cell size $p=5$ and $U=0.1$ for phase offsets $k=1$ and $k=2$. (Note that the soliton bands are almost flat with a bandwidth less than the width of the lines.) In the first case (Fig. 3 a, b) there is a single lowest-energy soliton solution with many-body Chern number, Eq. (11), $C=1$ and integer transport. In the second case (Fig. 3 c, d) two energies cross with a non-trivial Wilson loop, Eq. (18), $C_{2}=1$ giving rise to fractional transport of $1/2$. The data is calculated within the three-site soliton ansatz.
With the same approach we can numerically calculate the Berry curvature ${\cal F}(K,t)$ and from this the Chern number or Wilson loop. In Fig. 4 we have given an example for the Berry curvature of the AAH-model with $N=3$ particles, a $p=5$ unit-cell and a $k=1$ phase-offset, which shows integer quantized transport similar to Fig. 3. Using Eq. 11 for this Berry curvature yields a COM Chern number $C=1$. The yellow lines in Fig. 4 are numerical relicts and do not influence the result of the integration since they are effectively sets of measure zero.
Figure 4: Logarithm of the absolute value of the Berry curvature $\ln|{\cal F}(K,t)|$ in the lowest energy solution for $N=3$ particles, a $p=5$ unit-cell and a $k=1$ phase-offset. Integrating Eq. 11 with this Berry curvature yields a Chern number $C=1$. The sharp peaks, where $\ln|{\cal F}(K,t)|$ is on the order of 10, are numerical relicts but are effectively a set of measure zero for the integration.
We here showed the applicability of our COM approach to identify the different topological phases. In the following we will provide an explanation for the phase transitions.
@@ V Topological phase transitions
In the following we provide an explanation for the transitions between phases with integer, fractional and vanishing transport observed in the experiments.
V.1 Effective soliton Hamiltonian
The COM dynamics of bound $N$-particle objects (solitons) can be described by an effective single-particle Hamiltonian with eigenstates $|E_{\mu}(K)\rangle$:
@@ V.1 Effective soliton Hamiltonian
Since such a model misses out the continuum of extended states, it is only adequate for stable soliton solutions. Defining annihilation and creation operators for solitons centered at lattice site $l$ as $\hat{d}_{l}$ and $\hat{d}_{l}^{\dagger}$, respectively, the effective soliton Hamiltonian would read in coordinate space
@@
where the $J_{l,\textrm{eff}}$ are the effective hopping rates of the soliton and the $\epsilon_{l,\textrm{eff}}$ effective local energies. (Note that we here have assumed only nearest neighbor hopping.) All topological properties, including the topological classification according to the Altland-Zirnbauer scheme [1, 33], as well as all phase transitions of solitons are determined by this effective Hamilton.
In the limit of strong attractive interactions all soliton solutions have a localization length of a single lattice constant, i.e. all particles are located at the same lattice site. In this limit, we can explicitly derive the effective Hamiltonian from perturbation theory. Furthermore, it is obvious that in this limit the number of soliton bands is the same as the number of single-particle bands. We will explicitly construct $\mathcal{H}_{\textrm{eff}}$ for $N=3$ particles in the strong interaction limit in Sec. VI. Here we will first discuss some of its general properties.
@@
In the Aubry-Andre-Harper model, Eq. (1), any small attraction $U_{0}$ is sufficient to form a bound state (lattice soliton) with an energy below the continuum of extended $N$-particle states, if $N$ is large. Increasing the attractive interaction, the energy of these solitons is lowered and the bands deform. Moreover, excited soliton solutions can emerge. If the soliton bands have a non-trival Chern number or Wilson loop, (fractional) quantized topological transport can be observed [14, 13].
V.2 Merging of soliton bands
@@
The experiments in [14, 13] and DNLSE simulations showed that for very large but still finite values of the interaction, the topological transport stops altogether. This can be understood as follows: For $U\gg J$ the localization of all soliton solutions is reduced to a single lattice site. Thus the contribution of the local interaction to the energy, UN(N1)/2UN(N-1)/2, is the same for all solitons . Using a perturbative ansatz, one can show that moreover the contribution of the kinetic energy can be disregarded as it scales as $\sim J(\frac{J}{U})^{N-1}$ (see Appendix and Ref. [27]). Thus, the effective soliton Hamiltonian becomes approximately diagonal and is entirely determined by local energies $\epsilon_{l,\textrm{eff}}$, which for the AAH model result from virtual hopping processes of a single particle from the soliton site $l$ to an empty neighboring site (and back) and are in second-order perturbation theory proportional to $\epsilon_{l}\sim(J_{l}^{2}(t)+J_{l+1}^{2}(t))/U$. As a consequence, all soliton bands cross at different points of the pump cycle and the topological transport is given by the total Wilson loop of all COM bands, which is always zero for general reasons.
@@
Figure 5: Merging of soliton energies obtained from exact diagonalization with increasing interaction strength $U$ for $N=3$ particles, a $p=5$ unit-cell and $k=2$ phase-offset. The number of unit cells is $L=6$. A soliton originally prepared at $t=0$ in one of the two bands will remain in this band if the changes are adiabatic and the soliton energies do not touch (Figs. a) and b)). Once the energies touch in Dirac-like cones, the solutions switch bands at every crossing (red curve in Fig. c). The width of the lines is larger than the width of the soliton bands in $K$ space.
The two limiting cases show that as $U$ is increased, initially separated soliton solutions must approach each other and eventually merge. As shown in Fig. 5, obtained from exact diagonalization (ED) simulations of a small system (30 lattices sites), they do so by forming Dirac-like cones which touch at a critical interaction $U_{c}$. In this case the Chern numbers of the lowest two soliton bands are no longer adequate topological invariants, but one has to consider the Wilson loop. The topological transport can then become fractional, provided the soliton solutions which started at $t=0$ are energetically well separated and switch bands at the Dirac-like points. We now show that this is indeed the case:
@@
Close to the crossing points, which we assume to take place at $t=t_{0}$, the non-Abelian Berry connection $\mathbf{A}_{K}$ takes on the form
$\mathbf{A}_{K}=\left(\begin{array}[]{cc}a|\tau|&-\frac{ib}{\tau}\\ \frac{ib}{\tau}&-a|\tau|\end{array}\right),$ (26)
@@
with $\tau=t-t_{0}$. Here, we assumed that apart from a common energy offset $E_{0}=-E_{1}=a\,|\tau|$, and thus $\langle E_{0}|\partial_{t}{\cal H}|E_{1}\rangle/(E_{1}-E_{0})\approx b/\tau$. Due to the diverging off-diagonal elements, the eigensolutions no longer follow the original solutions but cross from one to the other (see color code in Fig. 5 c). Since $p$ is a prime number, there is an odd number of crossing points in one period in the AAH model. Thus, two periods are required for a soliton to return to the solution it started from. The shift in the COM position is then integer quantized only after two periods and the average transport per cycle is fractionally quantized. This integer-to-fractional transition is sharp, i.e. it does happen without a region of non-quantized transport in-between the two phases. The following section will however discuss a situation where an intermediate region of non-quantized transport appears.
V.3 Failure of transport quantization
@@
Based on solutions of the mean-field DNLSE it was shown in Ref. [42] for the example of the Rice-Mele model [32, 23, 9] with local attractive interactions that the transition between quantized topological pumping of a soliton to zero pumping upon increasing the interaction strength may go through an intermediate regime of non-quantized transport. This transient failure of transport quantization, where $\langle\Delta\hat{X}\rangle$ strongly fluctuates, was attributed in [42] to a ”self-crossing” of solutions of the semiclassical DNLSE. This phenomenon can be easily understood in the full quantum picture. For intermediate interactions, the excited soliton solution is in some parts of the pump cycle degenerate with the continuum of extended states and thus unstable. If this solution merges with the lower soliton band, the time evolution is no longer adiabatic and the transport is not quantized until the interaction is large enough such that also the excited band becomes fully gapped. In Fig. 6 we show the soliton energies in the Rice-Mele model with different values of the attractive onsite interaction $U$ obtained from exact diagonalization simulations of a $N=4$ soliton. The width of the lines reflects the width of the soliton bands $E(K,t)$ in $K$ space. Although only marginally visible due to finite size effects, Fig. 6b indicates the existence of a parameter range where the lowest soliton band becomes degenerate with an unstable excited solution leading to a fluctuating, non-quantized transport.
@@
Figure 6: Breakdown of transport quantization in the Rice-Mele model for intermediate values of onsite interactions. Shown are lowest energies for $N=4$ bosons. a) $U=2.0$: lowest soliton (quadron) band is gapped with Chern number $C=1$ ($\langle\Delta\hat{X}\rangle=1$). b) $U=3.0$: lowest two bands merge but (except for finite-size effects) overlap with extended states (unstable quadron) where the transport is not quantized. c) $U=4.0$ lowest two quadron bands merge with vanishing Wilson loop ($\langle\Delta\hat{X}\rangle=0$). The width of the lines reflects the width of the soliton bands in $K$ space.
VI Effective Hamiltonian: triplon model
@@
As stated in Sec. V.1, we can explicitly construct the effective soliton Hamiltonian in the strongly interacting limit where the non-interacting part of the Hamiltonian, Eq. (1), can be treated as a small perturbation and the solitons become maximally localized (i.e. with a localization length $\xi$ smaller than the lattice spacing).
@@
The doublon model ($N=2$) can also be treated analytically and, depending on the lattice model, can as well show an integer, fractional or trapped case. But, since in the doublon case both the effective hopping and the effective potential are of the same order in interaction strength $U^{-1}$, there is no qualitative change of the bandstructure and therefore no phase-transition.
Furthermore, while the dispersion of wavepackets during a topological pump can be suppressed also for single particles in lattices with weakly dispersive (or flat) bands [10], the interaction-driven topological phase transitions, discussed here, require virtual hopping processes which do not exists in single-particle systems. Thus these transitions are a remarkable and unique feature of composite particles (with more than two particles).
Figure 8: Instantaneous soliton energies for 3 particles for attractive interaction $U/J=1/3$ with phase-offset $k=8$ and unit-cell size $p=17$ with the corresponding COM-movement in a pump cycle of the red marked energy. The solutions are obtained with the three-site soliton ansatz.
VII Topological transport in higher soliton bands
So far we mainly discussed the behavior of ground-state solitons. In single-particle Thouless pumps (like the AAH-model without interaction) quantized topological transport also occurs in higher bands, determined by the corresponding Chern number, where the sum of all Chern-numbers has to vanish. While this is still true in the effective single-particle model, it does not cover the full physics in the composite particle setting. Here, excited composite-particle solutions might be unstable (compare Fig. 6), i.e. the soliton can be energetically degenerate with extended, i.e. not self-bound many-particle states. In such a case the state would not survive in the real time-evolution of the pump because the soliton can evaporate and therefore lose its composite structure.
If, on the other hand, an excited soliton is stable, there is quantized topological transport determined by the effective Chern number of the excited soliton band, see e.g. Fig. 1a in the extended data of Ref. [14]. In Fig. 8 we show the energy spectrum from the three-site soliton ansatz of the AAH model for $N=3$ particles with a $p=17$ unit-cell and a $k=8$ phase-offset. The interaction strength is choosen as $U/J=1/3$. As one can see, there are crossing solutions in the low-lying part of the spectrum, above which another gapped soliton solution exists (marked in red). The corresponding Chern number of this soliton band is $C=2$, which is in agreement with the displayed transport.
@@ VII Topological transport in higher soliton bands
We developed a full quantum description of topological pumps of self-bound $N$-particle states, i.e. lattice solitons. The quantum description allowed us to identify and to explicitly calculate a topological invariant, i.e. an effective single-particle Chern-number or Wilson-loop explaining the emergence of integer and fractional transport in the full range of interaction strength, where perturbative arguments [15, 28] fail. Transitions between phases of differently quantized topological transport observed in experiments [14, 13] as well as the possibility of parameter regimes without quantized transport were explained by coalescence of soliton bands and possible degeneracies with extended states. Quantized topological transport can also be observed in excited soliton bands, provided these bands are gapped and stable, i.e. there are no degeneracies with extended many-particle states. The concept can easily be extended to multi-component solitons [25, 38, 39], were, among other things, fractional transport was predicted despite the fact that all single-particle bands are topologically trivial. In the latter case interactions not only lead to the modification of topological properties, such as the transition from integer to fractional phases, but to the emergence of topology. Finally, we note that the concept can also be applied to bound many-particle states in cold gas experiments and topological pumps [43, 47, 4, 45, 44].
@@ Acknowledgements
Effective soliton Hamiltonians in the limit of strongly localized solitons